OpenStudy (anonymous):

Can someone tell me if I got the correct answer for this equation? x/(x-2) + (3x)/(x^2+7x+10) = (-3)/(x-5) I got x= 2, -3?

OpenStudy (anonymous):

\[\frac{x}{x-2}+\frac{3 x}{x^2+7 x+10}\text{=}\frac{-3}{x-5} \]The real roots of the above to 10 digit precision is:\[\{x\to -10.67296728,x\to 3.736912927\} \]